can you please show me the solution and answer to this problem, my daughter is having a hard time time on this one......five times a number diminished by 9 exceeds twice the number by 6. find the number
Answer by ptaylor(1894)

Let x=the number
2x=twice the number
2x+6=six more than twice the number (exceeds twice the number by 6)
5x=five times the number
5x-9=five times the number diminished by 9
So we have :
5x-9=2x+6 subtract 2x from and add 9 to both sides
5x-2x-9+9=2x-2x+6+9 collect like terms
3x=15
x=5-------------------the number we are looking for
CK
5x-9=2x+6
25-9=10+6
16=16


Hope this helps----ptaylor